The atomic planes in a graphite crystal are separated by 335 pm. At what angle would you find the first-order (n = 1) diffraction of 0.154 nm X-rays from a graphite crystal? Shows the calculation steps.
1
Expert's answer
2020-09-16T05:11:54-0400

In application of Braggs law

"2d.Sin\theta=n\lambda"

"Sin \theta=" "n\lambda\over 2d"

"\theta=unknown"

"n=1"

"\lambda=154pm"

"d=335pm"

"\therefore" "Sin \theta" "=" "1\times 154\over 2\times 335" "=0.229"

"Sin \theta =0.229"

"\theta= 13.24°"